OpenStudy (anonymous):

A solution of H2SO4(aq) with a molal concentration of 3.97 m has a density of 1.216 g/mL. What is the molar concentration of this solution?

OpenStudy (australopithecus):

I think you mean molality (m)? molality = Moles of Solute/Kg of Solution so we have, 3.97m = moles of H2SO4/ 1 Kg Note: I picked 1kg of solution arbitrarily, it doesn't matter what volume you pick you will still get the same molarity. solve for the moles of solute 3.97m * 1kg = moles of H2SO4 moles of H2SO4 = 3.97mol so now we need to convert 1kg to mL of solution 1kg = 1000g so set up the ratio, 1.216 g/mL = 1000g/xmL 1.216x = 1000 x = 1000/1.216 x= 822.37mL Volume of solution = 822.37mL 1mL = 1000L 822.37mL/1000 = 0.82237L Molarity = Moles of Solute (H2SO4)/Liters of Solution Molarity = 3.97mol/0.82237L Molarity = 4.83M

OpenStudy (australopithecus):

I wish I could edit posts, Error 1: Molality (m) = Moles of Solute/Kg of SOLVENT Error 2: "Note: I picked 1kg of solution arbitrarily, it doesn't matter what volume you pick you will still get the same molarity." should be, "Note: I picked 1kg of solvent arbitrarily, it doesn't matter what mass of solvent you pick you will still get the same molarity."

OpenStudy (australopithecus):

Mass of Solvent doesn't affect molarity in this case because moles is constant between both solutions as well as density is being used to convert the total Mass of Solvent to the total volume of Solution

OpenStudy (australopithecus):

In this case: Solute is the H2SO4 Solvent is just the water (assuming that it is dissolved in water)

OpenStudy (australopithecus):

If you have any questions feel free to ask

OpenStudy (anonymous):

If M =\[\frac{ mol }{ sol }\] or solute/ solution shouldn't we consider that molal concentration has the mass of solvent in the denominator whereas molar concentration has volume of solution in the denominator. Be sure to sum the masses of solute and solvent before applying the density.

OpenStudy (australopithecus):

yeah I made a mistake :S I'm rusty

OpenStudy (australopithecus):

but since you know that you should be able to determine the correct answer so meh

OpenStudy (anonymous):

I'm actually doing this online and its not accepting my answer either.

OpenStudy (australopithecus):

ok well let me try to correct for it

OpenStudy (australopithecus):

xg/98.078g/mol = 3.97mol 3.97mol*98.078g/mol =389.4g 389.4g + 1000 = 1389.4g 1.216 = 1389.4g/x 1.216x = 1389.4 x = 1389.4/1.216 x = 1142.598mL x = 1.142598mL 3.97mol/1.142598mL =3.475M

OpenStudy (australopithecus):

or 3.48M

OpenStudy (australopithecus):

I wrote mL instead of L by accident

OpenStudy (australopithecus):

is that correct?

OpenStudy (anonymous):

yea that's it!!! Thanks, sorry for the late response I was trying to figure it out also.

OpenStudy (australopithecus):

No thank you for informing me about my mistake originally, this site is great for brushing up on my solution calculations while I'm trapped in university :S

OpenStudy (australopithecus):

if you have any questions about my math please inform me and I can explain it to you